A student is measuring the average initial rate over the first 10 seconds of the reaction. They collect the following data:

t = 0 s   [A] = 0.131 mol L-1

t = 10 s  [A] = 0.097 mol L-1

What is the initial rate for this reaction? (Do not include units in your answer).



1
Expert's answer
2021-09-29T05:00:24-0400

0.131-0.097= 0.034

0.034/10 = 0.0034

= 3.4×10-3